# Exact computation of Transfer Entropy with Path Weight Sampling

### v1

@@ Appendix B TE-PWS algorithm
For clarity, a pseudocode to compute transfer entropy $\mathcal{T}_{X_{1}\to X_{3}}$ in only a three-variable process is given in Algorithm 1.
@@
1:inputs model parameters for simulating $\mathbf{X}(t)$
2:parameters timestep $\delta t$; number of timesteps $N$; number of trajectories for Monte-Carlo averages $M_{1}$ and $M_{2}$; resampling indicator $\kappa$
3:initialize Define trajectory labels $\nu$ and $\mu$; timestep variable $k$; cumulative transfer entropy array $\mathcal{T}[0:N]$; logarithm of path probabilities of $X_{3}$ within the spaces $(X_{1},X_{3})$ and $(X_{3})$ as $\mathcal{T}_{a}^{(\nu)}[0:N]$ and $\mathcal{T}_{b}^{(\nu)}[0:N]$ respectively.
4:$\nu\leftarrow 0$
@@
6:repeat
7:     Generate $M_{1}$ trajectories $(X_{1,[0,N]}^{(\nu)},X_{2,[0,N]}^{(\nu)},X_{3,[0,N]}^{(\nu)})$ jointly.
8:     $k\leftarrow 0$triangleright Accessing $P(X_{2,[0,N]}|X_{1,[0,N]},X_{3,[0,N]})$
9:     $\kappa\leftarrow M_{2}$
10:     initialize Generate $M_{2}$ samples of initial conditions $X_{2}^{(\mu)}(0)$ labeled by $\mu$ from a steady-state trajectory; weights in log scale $w^{(\mu)}=0$.
11:     repeat
12:         if $\kappa<M_{2}/2$ then
13:              Resample $M_{2}$ configurations from $X_{2}^{(\mu)}(k)$ with weights $\exp w^{(\mu)}$.
14:              For a jump process, additionally regenerate random values for the first waiting time and the first jump in each trajectory using the jump propensities.
@@
20:         Compute $\mathcal{T}_{a}^{(\nu)}[k]$ using Eq. 8 or 9 for a diffusion or jump process respectively.
@@
29:              For a jump process, additionally regenerate random values for the first waiting time and the first jump in each trajectory using the jump propensities.
30:              $w^{(\mu)}\leftarrow 0$ for all $\mu$
31:         end if
32:         Propagate reference dynamics $(X_{1,[k,k+1]}^{(\mu)},X_{2,[k,k+1]}^{(\mu)})$ as samples from $P_{0}(X_{1,[k,k+1]},X_{2,[k,k+1]}|X_{1,[0,k]},X_{2,[0,k]})$.
33:         $w^{(\mu)}\leftarrow w^{(\mu)}+\ln P\left(X_{1,[k,k+1]}^{(\mu)},X_{2,[k,k+1]}^{(\mu)},X_{3,[k,k+1]}^{(\nu)}\big{|}X_{1,[0,k]}^{(\mu)},X_{2,[0,k]}^{(\mu)},X_{3,[0,k]}^{(\nu)}\right)$
34:         $w^{(\mu)}\leftarrow w^{(\mu)}-\ln P_{0}\left(X_{1,[k,k+1]}^{(\mu)},X_{2,[k,k+1]}^{(\mu)}\big{|}X_{1,[0,k]}^{(\mu)},X_{2,[0,k]}^{(\mu)}\right)$
35:         $\kappa\leftarrow(\sum_{\mu}\exp w^{(\mu)})^{2}/\sum_{\mu}\exp(2w^{(\mu)})$
36:         Compute $\mathcal{T}_{b}^{(\nu)}[k]$ using Eq. 7 or 10 for a diffusion or jump process respectively.
37:         $k\leftarrow k+1$
38:     until $k=N$
39:     $\mathcal{T}[0:N]\leftarrow\mathcal{T}[0:N]+\mathcal{T}^{(\nu)}_{a}[0:N]-\mathcal{T}^{(\nu)}_{b}[0:N]$
40:     $\nu\leftarrow\nu+1$
41:until $\nu=M_{1}$
@@ Appendix C Proof of the efficiency of the RR scheme
TE-PWS performs the average in Eq. 7 with high statistical efficiency by preferentially sampling rare large values of the summand. As discussed in [13], this is implemented with a stochastic particle filter through which the ensemble of $M_{2}$ trajectories is resampled after every $\delta t$ time. For example, in order to compute $P(X_{j,[0,N]}^{(\nu)})$, an ensemble of $X_{i}$ and $X_{l}$ trajectories is simulated. At the $(k+1)$-th step, the resampling weight used for the $\mu$-th trajectory in the ensemble is
$\widehat{g}\big{[}X_{l,[0,k+1]}^{(\mu)}\big{]}$
$=P\left(X_{i,[k,k+1]}^{(\mu)},X_{j,[k,k+1]}^{(\nu)},X_{l,[k,k+1]}^{(\mu)}\Big{|}X_{i,[0,k]}^{(\mu)},X_{j,[0,k]}^{(\nu)},X_{l,[0,k]}^{(\mu)}\right)$
$\qquad/\;P_{0}\left(X_{i,[k,k+1]}^{(\mu)},X_{l,[k,k+1]}^{(\mu)}\Big{|}X_{i,[0,k]}^{(\mu)},X_{l,[0,k]}^{(\mu)}\right)$ (15)
@@
which is analytically available. Using this weight, we show below that iterative resampling changes the trajectory distribution optimally such that the summand in Eq. 7 is a constant, i.e., has a variance of zero, achieving perfect sampling. We will show this by first showing that if after the $k$-th step the trajectories have an optimal distribution, they stay optimal after the $(k+1)$-th step. Then, combined with the fact that the distribution of initial conditions is by construction optimal, we will conclude by induction that the distribution stays uniform during the entire duration of the trajectory.
@@
The optimal choice for $P_{0}$ would be the hypothetical $P_{0}(X_{i,[0,N]}^{(\mu)},X_{l,[0,N]}^{(\mu)})=P(X_{i,[0,N]}^{(\mu)},X_{l,[0,N]}^{(\mu)}|X_{j,[0,N]}^{(\nu)})$ as it makes the summand in Eq. 7 a constant independent of the index $\mu$. The purpose of the particle filter is to bias the simulated distribution $P_{0}(X_{i,[0,N]}^{(\mu)},X_{l,[0,N]}^{(\mu)})$ towards the optimal distribution $P(X_{i,[0,N]}^{(\mu)},X_{l,[0,N]}^{(\mu)}|X_{j,[0,N]}^{(\nu)})$. For the inductive argument, assume that after the $k$-th step, the trajectories are distributed optimally according to $P(X_{i,[0,k]}^{(\mu)},X_{l,[0,k]}^{(\mu)}|X_{j,[0,k]}^{(\nu)})$. After the propagation step, the probability of each trajectory changes to a product of its previous value and the probability of the new segment,
$w^{(\mu)}$ $=P\left(X_{i,[0,k]}^{(\mu)},X_{l,[0,k]}^{(\mu)}\Big{|}X_{j,[0,k]}^{(\nu)}\right)$
.$\qquad.\;P_{0}\left(X_{i,[k,k+1]}^{(\mu)},X_{l,[k,k+1]}^{(\mu)}\Big{|}X_{i,[0,k]}^{(\mu)},X_{l,[0,k]}^{(\mu)}\right)$
$=P\left(X_{i,[0,k]}^{(\mu)},X_{j,[0,k]}^{(\nu)},X_{l,[0,k]}^{(\mu)}\right)$
.$\qquad.\;P_{0}\left(X_{i,[k,k+1]}^{(\mu)},X_{l,[k,k+1]}^{(\mu)}\Big{|}X_{i,[0,k]}^{(\mu)},X_{l,[0,k]}^{(\mu)}\right)$
$\qquad/\;P\left(X_{j,[0,k]}^{(\nu)}\right)$ (16)
@@
Then we resample the trajectories with weight $\widehat{g}$ given in Eq. 15. The probability after resampling becomes proportional to the product of the Eqs. 15 and 16,
$w^{(\mu)}\widehat{g}\big{[}X_{l,[0,k+1]}^{(\mu)}\big{]}$ $=\frac{P\left(X_{i,[0,k+1]}^{(\mu)},X_{j,[0,k+1]}^{(\nu)},X_{l,[0,k+1]}^{(\mu)}\right)}{P\left(X_{j,[0,k]}^{(\nu)}\right)}$ (17)
@@
where Bayes’ theorem has been used to condense the numerator. The normalization constant for this probability is obtained by summing over all $X_{i,[0,k+1]}^{(\mu)}$ and $X_{l,[0,k+1]}^{(\mu)}$ trajectories, which gives $P\left(X_{j,[0,k+1]}^{(\nu)}\right)\big{/}P\left(X_{j,[0,k]}^{(\nu)}\right)$. Dividing Eq. 17 by this normalization constant gives the new normalized probability distribution as $P(X_{i,[0,k+1]}^{(\mu)},X_{l,[0,k+1]}^{(\mu)}|X_{j,[0,k+1]}^{(\nu)})$.
Hence the trajectories remain distributed optimally after resampling. Additionally, initial conditions for the $M_{2}$ trajectories are sampled from the same distribution as the original $M_{1}$ trajectories. Then by induction, the resampling procedure, i.e., the RR scheme, gives access to the exact conditional distribution at every step. Resampling is algorithmically performed with a stratified resampling technique which is computationally efficient32. The optimal distribution is then used in Eq. 7 to compute the denominator in Eq. 5, a part of the transfer entropy. The numerator is obtained similarly by accessing the corresponding conditional distribution with the RR scheme. Conditional distributions in trajectory space for all combinations of variables can be accessed by repeating this procedure.
@@ Appendix D Choice of reference distribution
The transfer entropy estimated from TE-PWS is exact for any choice of the reference probability $P_{0}$ owing to the RR scheme. However, the number of times pruning and enrichment needs to be performed depends on how large the variance of the summand in Eq. 7 is. A better choice of $P_{0}$ results in a smaller variance at the same computational cost. As discussed earlier, for computing $P(X_{j,[0,N]}^{(\nu)})$ for example, the ideal choice for $P_{0}$ would be $P_{0}(X_{i,[0,N]}^{(\mu)},X_{l,[0,N]}^{(\mu)})=P(X_{i,[0,N]}^{(\mu)},X_{l,[0,N]}^{(\mu)}|X_{j,[0,N]}^{(\nu)})$, which is not known a priori and is impossible to directly sample from. We instead choose a distribution $P_{0}(X_{i,[0,N]}^{(\mu)},X_{l,[0,N]}^{(\mu)})$ that uses the past trajectory of $X_{j,[0,N]}^{(\nu)}$ at every timestep to compute the drift and diffusion terms, similar to the original dynamics of the system in the full $d$-dimensional space. This keeps the reference distribution close to the target conditional distribution while being analytically known and easy to sample from.
@@
As an example, consider the three-dimensional OU process discussed before, $\dot{\mathbf{X}}=-\mathbf{a}\mathbf{X}+\bm{\xi}$ where $\bm{\xi}$ is a Gaussian white noise with zero mean and covariance matrix $2\mathbf{D}$. The diffusion constant matrix $\mathbf{D}$ may have nonzero off-diagonal elements. In the full three-dimensional space, propagating the natural dynamics involves computing the drifts and sampling three correlated noise components from the joint Gaussian distribution, which we will call $\mathbb{G}(\xi_{i},\xi_{j},\xi_{l})$. For sampling from a $P_{0}(X_{i,[0,N]}^{(\mu)},X_{l,[0,N]}^{(\mu)})$ that is as close as possible to $P(X_{i,[0,N]}^{(\mu)},X_{l,[0,N]}^{(\mu)}|X_{j,[0,N]}^{(\nu)})$, we propagate $X_{i}^{(\mu)}$ and $X_{l}^{(\mu)}$ trajectories with the equations of motion
$\dot{X}_{i}^{(\mu)}(k)$ $=-a_{11}X_{i}^{(\mu)}(k)-a_{12}X^{(\nu)}_{j}(k)-a_{13}X_{l}^{(\mu)}(k)$
$+\tilde{\xi}_{i}^{(\mu)}(k)$ (18)
$\dot{X}_{l}^{(\mu)}(k)$ $=-a_{31}X_{i}^{(\mu)}(k)-a_{32}X^{(\nu)}_{j}(k)-a_{33}X_{l}^{(\mu)}(k)$
$+\tilde{\xi}_{l}^{(\mu)}(k)$ (19)
@@
where $\tilde{\xi}_{i}^{(\mu)}$ and $\tilde{\xi}_{l}^{(\mu)}$ are Gaussian white noises whose distribution should be commensurate with the existing noises in the $X_{j,[0,N]}^{(\nu)}$ trajectory. However, given only the $X_{j,[0,N]}^{(\nu)}$ trajectory, the noises $\xi_{j}^{(\nu)}(k)$ are not uniquely known, as equivalently the drifts $F_{j}^{(\nu)}(k)$ are unknown. We now make an additional choice of deriving approximate noises for $X_{j,[0,N]}^{(\nu)}$, called $\tilde{\xi}_{j}^{\nu\mu}$, by assuming that the drift in $X_{j}$ was derived in turn from the $X_{i,[0,k]}^{(\mu)}$ and $X_{j,[0,k]}^{(\mu)}$ trajectories. Using the given $X_{j}^{(\nu)}$ trajectory, at every timestep, we solve
$\dot{X}_{j}^{(\nu)}(k)$ $=-a_{21}X_{i}^{(\mu)}(k)-a_{22}X^{(\nu)}_{j}(k)-a_{23}X_{l}^{(\mu)}(k)$
$+\tilde{\xi}_{j}^{(\nu\mu)}(k)$ (20)
@@
for $\tilde{\xi}_{j}^{(\nu\mu)}(k)$, which can be done by simply transposing the equation even if the original dynamics was not linear. Now given $\tilde{\xi}_{j}^{(\nu\mu)}(k)$, from which distribution should we generate $\tilde{\xi}_{i}^{(\mu)}(k)$ and $\tilde{\xi}_{l}^{(\mu)}(k)$? Here we use the conditional noise distribution in the natural dynamics of the system in order to keep the reference distribution close to optimal. Given the multivariate Gaussian distribution $\mathbb{G}(\xi_{i},\xi_{j},\xi_{l})$ for the original system of noises, and given that we have specified $\xi_{j}=\tilde{\xi}_{j}^{(\nu\mu)}(k)$, the distribution we need to sample from is given by the conditional distribution $\mathbb{G}(\xi_{i}=\tilde{\xi}_{i},\xi_{l}=\tilde{\xi}_{l}|\xi_{j}=\tilde{\xi}_{j}^{(\nu\mu)}(k))$. Using the Schur complement formula, the mean and the covariance matrix of this distribution are given by $\Sigma_{0}\Sigma^{-1}\tilde{\xi}_{j}^{(\nu\mu)}$ and covariance matrix $2\widetilde{D}-\Sigma_{0}\Sigma^{-1}\Sigma_{0}^{T}$ respectively, where
$\Sigma_{0}$ $=2\begin{pmatrix}D_{ij}\\ D_{lj}\end{pmatrix},$ (21)
$\Sigma$ $=2D_{jj},\mathrm{~{}and}$ (22)
$\widetilde{D}$ $=\begin{pmatrix}D_{ii}&&D_{il}\\ D_{il}&&D_{ll}\end{pmatrix}.$ (23)
@@
The summary of the procedure for generating $X_{i}^{(\mu)}$ and $X_{l}^{(\mu)}$ trajectories for a given $X_{j}^{(\nu)}$ trajectory consists of the following steps at every timestep.
  • The drifts for propagating $X_{i}^{(\mu)}(k)$ and $X_{l}^{(\mu)}(k)$ are obtained using $X_{j}^{(\nu)}(k)$.

### v2

@@ Appendix B Pseudocode
A pseudocode for computing the transfer entropy $\mathcal{T}_{X_{1}\to X_{3}}$ in a three-variable stochastic process via the TE-PWS algorithm is provided in Algorithm 1.
@@
1:objective computation of the transfer entropy $\mathcal{T}_{X_{1}\to X_{3}}$ in a three-variable stochastic process.
2:inputs model parameters for simulating $\mathbf{X}(t)$
3:parameters timestep $\delta t$; number of timesteps $N$; number of trajectories for Monte-Carlo averages $M_{1}$ and $M_{2}$; resampling indicator $\kappa$
4:initialize Define trajectory labels $\nu$ and $\mu$; timestep variable $k$; cumulative transfer entropy array $\mathcal{T}[0:N]$; two terms (factors) of transfer entropy $\mathcal{T}_{a}^{(\nu)}[0:N]$ and $\mathcal{T}_{b}^{(\nu)}[0:N]$ respectively.
5:$\nu\leftarrow 0$
@@
7:repeat
8:     Generate $M_{1}$ trajectories $(X_{1,[0,N]}^{(\nu)},X_{2,[0,N]}^{(\nu)},X_{3,[0,N]}^{(\nu)})$ jointly.
9:     $k\leftarrow 0$triangleright Accessing $P(X_{2,[0,N]}|X_{1,[0,N]},X_{3,[0,N]})$
10:     $\kappa\leftarrow M_{2}$
11:     initialize Generate $M_{2}$ samples of initial conditions $X_{2}^{(\mu)}(0)$ labeled by $\mu$ from a steady-state trajectory; weights in log scale $w^{(\mu)}=0$.
12:     repeat
13:         if $\kappa<M_{2}/2$ then
14:              Resample $M_{2}$ configurations from $X_{2}^{(\mu)}(k)$ with weights $\exp w^{(\mu)}$.
@@
20:         Compute $\mathcal{T}_{a}^{(\nu)}[k]$ using Eq. 8 or 9 of the main text for a diffusion or jump process respectively.
@@
29:              $w^{(\mu)}\leftarrow 0$ for all $\mu$
30:         end if
31:         Propagate reference dynamics $(X_{1,[k,k+1]}^{(\mu)},X_{2,[k,k+1]}^{(\mu)})$ as samples from $P_{0}(X_{1,[k,k+1]},X_{2,[k,k+1]}|X_{1,[0,k]},X_{2,[0,k]})$.
32:         $w^{(\mu)}\leftarrow w^{(\mu)}+\ln P\left(X_{1,[k,k+1]}^{(\mu)},X_{2,[k,k+1]}^{(\mu)},X_{3,[k,k+1]}^{(\nu)}\big{|}X_{1,[0,k]}^{(\mu)},X_{2,[0,k]}^{(\mu)},X_{3,[0,k]}^{(\nu)}\right)$
33:         $w^{(\mu)}\leftarrow w^{(\mu)}-\ln P_{0}\left(X_{1,[k,k+1]}^{(\mu)},X_{2,[k,k+1]}^{(\mu)}\big{|}X_{1,[0,k]}^{(\mu)},X_{2,[0,k]}^{(\mu)}\right)$
34:         $\kappa\leftarrow(\sum_{\mu}\exp w^{(\mu)})^{2}/\sum_{\mu}\exp(2w^{(\mu)})$
35:         Compute $\mathcal{T}_{b}^{(\nu)}[k]$ using Eq. 7 or 10 in the main text for a diffusion or jump process respectively.
36:         $k\leftarrow k+1$
37:     until $k=N$
38:     $\mathcal{T}[0:N]\leftarrow\mathcal{T}[0:N]+\mathcal{T}^{(\nu)}_{a}[0:N]-\mathcal{T}^{(\nu)}_{b}[0:N]$
39:     $\nu\leftarrow\nu+1$
40:until $\nu=M_{1}$
@@ Appendix C Proof of the efficiency of the RR scheme
Here we prove that the Rosenbluth-Rosenbluth (RR) scheme in TE-PWS achieves optimal efficiency for a Monte-Carlo evaluation of the transfer entropy. For clarity, we start by reproducing Eq. 7 of the main text here, representing a typical marginalization procedure in TE-PWS,
$P\left(X_{j,[0,k]}^{(\nu)}\right)$ $=\frac{1}{M_{2}}\sum_{\mu}\frac{P\left(X_{i,[0,k]}^{(\mu)},X_{j,[0,k]}^{(\nu)},X_{l,[0,k]}^{(\mu)}\right)}{P_{0}\left(X_{i,[0,k]}^{(\mu)},X_{l,[0,k]}^{(\mu)}\right)}$ (S1)
@@
TE-PWS performs this average with high statistical efficiency by preferentially sampling rare large values of the summand. Similar to PWS as discussed in [1], this is implemented using the RR scheme, in which the ensemble of $M_{2}$ trajectories is resampled after every $\delta t$ time (see Fig. 1 of main text). For example, in order to compute $P(X_{j,[0,N]}^{(\nu)})$, an ensemble of $X_{i}$ and $X_{l}$ trajectories is simulated. At the $(k+1)$-th step, the resampling weight used for the $\mu$-th trajectory in the ensemble is
$\widehat{g}\big{[}X_{l,[0,k+1]}^{(\mu)}\big{]}=P\left(X_{i,[k,k+1]}^{(\mu)},X_{j,[k,k+1]}^{(\nu)},X_{l,[k,k+1]}^{(\mu)}\Big{|}X_{i,[0,k]}^{(\mu)},X_{j,[0,k]}^{(\nu)},X_{l,[0,k]}^{(\mu)}\right)/\;P_{0}\left(X_{i,[k,k+1]}^{(\mu)},X_{l,[k,k+1]}^{(\mu)}\Big{|}X_{i,[0,k]}^{(\mu)},X_{l,[0,k]}^{(\mu)}\right)$ (S2)
@@
which is analytically available. We show below that iteratively resampling the newly generated trajectories with this weight changes the trajectory distribution optimally such that the summand in Eq. S1 is a constant, i.e., has a variance of zero, achieving perfect sampling. We will show this by first showing that if after the $k$-th step the trajectories have an optimal distribution, they stay optimal after the $(k+1)$-th step. Then, combined with the fact that the distribution of initial conditions is by construction optimal, we will conclude by induction that the distribution stays uniform during the entire duration of the trajectory.
@@
The optimal choice for $P_{0}$ would be the hypothetical $P_{0}(X_{i,[0,N]},X_{l,[0,N]})=P(X_{i,[0,N]},X_{l,[0,N]}|X_{j,[0,N]}^{(\nu)})$ as it makes the summand in Eq. S1 a constant independent of the index $\mu$, resulting in a zero-variance estimate of $P(X_{j,[0,N]}^{(\nu)})$. The purpose of the RR scheme is to bias the simulated distribution $P_{0}(X_{i,[0,N]}^{(\mu)},X_{l,[0,N]}^{(\mu)})$ towards the optimal distribution $P(X_{i,[0,N]}^{(\mu)},X_{l,[0,N]}^{(\mu)}|X_{j,[0,N]}^{(\nu)})$. For the inductive argument, assume that after the $k$-th step, the trajectories are distributed optimally according to $P(X_{i,[0,k]}^{(\mu)},X_{l,[0,k]}^{(\mu)}|X_{j,[0,k]}^{(\nu)})$. After the next propagation step, the probability of each trajectory changes to a product of its previous value and the probability of the new segment,
$w^{(\mu)}$ $=P\left(X_{i,[0,k]}^{(\mu)},X_{l,[0,k]}^{(\mu)}\Big{|}X_{j,[0,k]}^{(\nu)}\right).\;P_{0}\left(X_{i,[k,k+1]}^{(\mu)},X_{l,[k,k+1]}^{(\mu)}\Big{|}X_{i,[0,k]}^{(\mu)},X_{l,[0,k]}^{(\mu)}\right)$
$=P\left(X_{i,[0,k]}^{(\mu)},X_{j,[0,k]}^{(\nu)},X_{l,[0,k]}^{(\mu)}\right).\;P_{0}\left(X_{i,[k,k+1]}^{(\mu)},X_{l,[k,k+1]}^{(\mu)}\Big{|}X_{i,[0,k]}^{(\mu)},X_{l,[0,k]}^{(\mu)}\right)\big{/}\;P\left(X_{j,[0,k]}^{(\nu)}\right)$ (S3)
@@
Then we resample the trajectories with weight $\widehat{g}$ given in Eq. S2. The probability distribution of these trajectories after this resampling step becomes proportional to the product of Eqs. S2 and S3,
$w^{(\mu)}\widehat{g}\big{[}X_{l,[0,k+1]}^{(\mu)}\big{]}$ $=\frac{P\left(X_{i,[0,k+1]}^{(\mu)},X_{j,[0,k+1]}^{(\nu)},X_{l,[0,k+1]}^{(\mu)}\right)}{P\left(X_{j,[0,k]}^{(\nu)}\right)}$ (S4)
@@
where Bayes’ theorem has been used to condense the numerator. The normalization constant for this probability is obtained by summing Eq. S4 over all $X_{i,[0,k+1]}^{(\mu)}$ and $X_{l,[0,k+1]}^{(\mu)}$ trajectories, which gives $P\left(X_{j,[0,k+1]}^{(\nu)}\right)\big{/}P\left(X_{j,[0,k]}^{(\nu)}\right)$. Dividing Eq. S4 by this normalization constant gives the new normalized probability distribution as $P(X_{i,[0,k+1]}^{(\mu)},X_{l,[0,k+1]}^{(\mu)}|X_{j,[0,k+1]}^{(\nu)})$. Hence the trajectories remain distributed optimally after resampling.
To complete the proof by induction, we also need to show that the initial conditions of the $X_{i,[0,k]}^{(\mu)}$ and $X_{l,[0,k]}^{(\mu)}$ trajectories are consistent with the optimal distribution $P(X_{i,[0,k]}^{(\mu)},X_{l,[0,k]}^{(\mu)}|X_{j,[0,k]}^{(\nu)})$. This requirement is met by construction, because the initial conditions for the $X_{i}^{(\mu)}$ and $X_{l}^{(\mu)}$ trajectories are sampled from the same joint distribution $P(\mathbf{X}(0))$ as that from which $X_{j}^{(\nu)}(0)$ is drawn; if one is interested in the steady-state transfer entropy rates, this distribution is the steady-state joint distribution where $X_{j}=X_{j}^{(\nu)}(0)$ is given. This ingredient, together with the fact that the resampling procedure preserves the optimal distribution as described above, guarantees that, by induction, the RR scheme generates samples from the optimal conditional distribution $P(X_{i,[0,k]},X_{l,[0,k]}|X_{j,[0,k]}^{(\nu)})$ at every step. Resampling is algorithmically performed with a stratified resampling technique which is computationally efficient2. Samples from the optimal distribution are then used in Eq. S1 to compute the denominator in Eq. 5 of the main text, reproduced here,
$\mathcal{T}_{X_{i}\to X_{j}}$ $=\frac{1}{M_{1}}\sum_{\nu}\sum_{k}\ln\frac{P\left(X^{(\nu)}_{j}(k+1)\big{|}X^{(\nu)}_{i,[0,k]},X^{(\nu)}_{j,[0,k]}\right)}{P\left(X^{(\nu)}_{j}(k+1)\big{|}X^{(\nu)}_{j,[0,k]}\right)}$ (S5)
The numerator is obtained similarly by generating samples from the corresponding conditional distribution $P(X_{l,[0,k]}|X_{i,[0,k]}^{(\nu)},X_{j,[0,k]}^{(\nu)})$ with the RR scheme. Conditional distributions of trajectories of all combinations of variables can be similarly sampled by repeating this procedure.
@@ Appendix D Choice of reference distribution
The transfer entropy estimated from TE-PWS is exact for any choice of the reference probability $P_{0}$ because of the RR scheme. However, the number of times resampling needs to be performed depends on how large the variance of the summand in Eq. S1 is. A better choice of $P_{0}$ results in a smaller variance at the same computational cost. As discussed earlier, for computing $P(X_{j,[0,N]}^{(\nu)})$ for example, the ideal choice for $P_{0}$ would be $P_{0}(X_{i,[0,N]}^{(\mu)},X_{l,[0,N]}^{(\mu)})=P(X_{i,[0,N]}^{(\mu)},X_{l,[0,N]}^{(\mu)}|X_{j,[0,N]}^{(\nu)})$, which is not known a priori and is impossible to directly sample from. We therefore choose a distribution $P_{0}(X_{i,[0,N]}^{(\mu)},X_{l,[0,N]}^{(\mu)})$ that uses the past trajectory of $X_{j,[0,N]}^{(\nu)}$ at every timestep to compute the drift and diffusion terms for $X_{i}$ and $X_{l}$, similar to the original dynamics of the system in the full $d$-dimensional space. This keeps the reference distribution close to the target conditional distribution while being analytically known and easy to sample from.
@@
As an example, consider the three-dimensional OU process discussed before, $\dot{\mathbf{X}}=-\mathbf{a}\mathbf{X}+\bm{\xi}$ where $\mathbf{a}$ is the spring-constant matrix and $\bm{\xi}$ is a Gaussian white noise with zero mean and covariance matrix $2\mathbf{D}$. The diffusion constant matrix $\mathbf{D}$ may have nonzero off-diagonal elements. In the full three-dimensional space, propagating the natural dynamics involves computing the drifts and sampling three correlated noise components from the joint Gaussian distribution, which we will call $\mathbb{G}(\xi_{i},\xi_{j},\xi_{l})$. For sampling trajectories $X_{i,[0,N]}^{(\mu)}$ and $X_{l,[0,N]}^{(\mu)}$, for a given $X_{j,[0,N]}^{(\nu)}$, from a distribution $P_{0}(X_{i,[0,N]}^{(\mu)},X_{l,[0,N]}^{(\mu)})$ that is as close as possible to $P(X_{i,[0,N]}^{(\mu)},X_{l,[0,N]}^{(\mu)}|X_{j,[0,N]}^{(\nu)})$, we propagate $X_{i}^{(\mu)}$ and $X_{l}^{(\mu)}$ trajectories with the equations of motion
$\dot{X}_{i}^{(\mu)}(k)$ $=-a_{11}X_{i}^{(\mu)}(k)-a_{12}X^{(\nu)}_{j}(k)-a_{13}X_{l}^{(\mu)}(k)+\tilde{\xi}_{i}^{(\mu)}(k)$ (S6)
$\dot{X}_{l}^{(\mu)}(k)$ $=-a_{31}X_{i}^{(\mu)}(k)-a_{32}X^{(\nu)}_{j}(k)-a_{33}X_{l}^{(\mu)}(k)+\tilde{\xi}_{l}^{(\mu)}(k)$ (S7)
@@
where $\tilde{\xi}_{i}^{(\mu)}$ and $\tilde{\xi}_{l}^{(\mu)}$ are Gaussian white noises whose distributions should be commensurate with the existing noises $\xi_{j}^{(\nu)}$ in the $X_{j,[0,N]}^{(\nu)}$ trajectory. However, given only the $X_{j,[0,N]}^{(\nu)}$ trajectory, the noises $\xi_{j}^{(\nu)}(k)$ are not uniquely known since the drifts $F_{j}^{(\nu)}(k)$ are unknown; hence, the distributions of $\tilde{\xi}_{i}^{(\mu)}$ and $\tilde{\xi}_{l}^{(\mu)}$ are not yet uniquely defined. To fully specify these distributions, we now make a choice for deriving approximate noises in $X_{j,[0,N]}^{(\nu)}$, called $\tilde{\xi}_{j}^{(\nu\mu)}$, by assuming that the drift in $X_{j}$ depends, besides $X_{j}^{(\nu)}(k)$ itself, on $X_{i}^{(\mu)}(k)$ and $X_{l}^{(\mu)}(k)$, i.e. the drift is taken to be $-a_{21}X_{i}^{(\mu)}(k)-a_{22}X^{(\nu)}_{j}(k)-a_{23}X_{l}^{(\mu)}(k)$. Next, obtaining $\dot{X}_{j}^{(\nu)}(k)$ from the given $X_{j}^{(\nu)}$ trajectory and with the drift now specified, we solve, at every timestep,
$\dot{X}_{j}^{(\nu)}(k)$ $=-a_{21}X_{i}^{(\mu)}(k)-a_{22}X^{(\nu)}_{j}(k)-a_{23}X_{l}^{(\mu)}(k)+\tilde{\xi}_{j}^{(\nu\mu)}(k)$ (S8)
@@
for $\tilde{\xi}_{j}^{(\nu\mu)}(k)$, which can be done by simply transposing the equation even if the original dynamics was not linear. Now given $\tilde{\xi}_{j}^{(\nu\mu)}(k)$, from which distribution should we generate $\tilde{\xi}_{i}^{(\mu)}(k)$ and $\tilde{\xi}_{l}^{(\mu)}(k)$? Here we use the conditional noise distribution in the natural dynamics of the system in order to keep the reference distribution close to optimal. Given the multivariate Gaussian distribution $\mathbb{G}(\xi_{i},\xi_{j},\xi_{l})$ for the original system of noises, and given that we have specified $\xi_{j}=\tilde{\xi}_{j}^{(\nu\mu)}(k)$, the distribution we need to sample from is given by the conditional distribution $\mathbb{G}(\xi_{i}=\tilde{\xi}_{i},\xi_{l}=\tilde{\xi}_{l}|\xi_{j}=\tilde{\xi}_{j}^{(\nu\mu)}$. Using the Schur complement formula, the mean and the covariance matrix of this distribution are given by $\Sigma_{0}\Sigma^{-1}\tilde{\xi}_{j}^{(\nu\mu)}$ and $2\widetilde{D}-\Sigma_{0}\Sigma^{-1}\Sigma_{0}^{T}$ respectively, where
$\Sigma_{0}$ $=2\begin{pmatrix}D_{ij}\\ D_{lj}\end{pmatrix},$ (S9)
$\Sigma$ $=2D_{jj},\mathrm{~{}and}$ (S10)
$\widetilde{D}$ $=\begin{pmatrix}D_{ii}&&D_{il}\\ D_{il}&&D_{ll}\end{pmatrix}.$ (S11)
@@
Summarizing, the procedure for generating $X_{i}^{(\mu)}$ and $X_{l}^{(\mu)}$ trajectories for a given $X_{j}^{(\nu)}$ trajectory consists of the following steps at every timestep:
  • The drifts for propagating $X_{i}^{(\mu)}(k)$ and $X_{l}^{(\mu)}(k)$ are obtained using $X_{i}^{(\mu)}(k)$, $X_{j}^{(\nu)}(k)$ and $X_{l}^{(\mu)}(k)$.